# PAS: A Training-Free Stabilizer for Temporal Encoding in Video LLMs

### v1

@@ Abstract
Video LLMs suffer from temporal inconsistency: small shifts in frame timing can flip attention and suppress relevant frames. We trace this instability to the common extension of Rotary Position Embeddings to video through multimodal RoPE. The induced inverse Fourier time kernel exhibits frame-scale ripples that multiply adjacent frames by different factors, which perturbs attention that should otherwise be governed by the raw query key inner product. We present Phase Aggregated Smoothing (PAS), a simple, training-free mechanism that applies small opposed phase offsets across heads and then aggregates their outputs. PAS preserves the per-head spectrum magnitude, while the aggregation effectively smooths the temporal kernel and reduces phase sensitivity without changing the positional encoding structure. Our analysis shows that the RoPE rotated logit can be approximated as a content dot product scaled by a time kernel; smoothing this kernel yields Lipschitz stability of attention to small temporal shifts; multi phase averaging attenuates high frequency ripples while preserving per-head spectra under Nyquist-valid sampling. Experiments on multiple video understanding benchmarks under matched token budgets show consistent improvements with negligible computational overhead. PAS provides a plug and play upgrade for robust temporal encoding in Video LLMs.
@@ 1 Introduction
Large multimodal language models extend language-centric architectures to video, enabling open-ended description and reasoning over long temporal contexts. A widely used positional encoding is Rotary Position Embeddings (RoPE), originally designed for discrete text tokens. In Video LLMs, RoPE is generalized to a multimodal variant (M-RoPE) that encodes temporal ($T$), height ($H$), and width ($W$) dimensions. Along the temporal axis, this construction amounts to assigning a finite set of temporal frequencies; each frequency line rotates query and key pairs and induces an independent, periodic contribution to the attention logit. In practice, downstream accuracy is sensitive to sampling details: small changes of frame rate or sampling offsets can alter which motion cues are aggregated per token, exposing fragility in temporal encoding. One of the reasons is that when this text-oriented mechanism is applied to continuous time, the attention distribution may be be disrupted.
Figure 1: A failure from a real clip [yt:CJHP6dPjuGY]. Temporal RoPE imposes an interval dependent gain on attention; when a key frame lies in a low gain trough of the modulation, it is down weighted compared to less relevant frames, which propagates to downstream errors.
@@
To make this precise, we adopt a Fourier view. Let $\{\omega_{k}\}$ denote the temporal line set and define $m(t)$ as the inverse Fourier time-domain kernel associated with these lines. Under a standard high dimensional approximation, the RoPE rotated logit at relative lag $t$ can be written as the unrotated content dot product multiplied by $\mathrm{Re}\{m(t)\}$; we later formalize this phase modulation view as Theorem 1. Because $m(t)$ is an average of cosines over $\{\omega_{k}\}$, it is intrinsically rippled at the frame scale, which renders the overall modulation non-smooth. Such non-smoothness means that the attention assigned to frames varies with their relative lag: within specific temporal intervals, the modulation can either amplify or attenuate attention. A concrete failure occurs when a key frame lands in a low gain interval of $m(t)$, causing its evidence to be suppressed or even ignored (figure 1). The upper panel of figure 2 visualizes how this interval dependent modulation undermines temporal stability by imposing gain swings across adjacent frames.
@@
    We identify and formalize a core limitation of applying text oriented RoPE to the time axis of video. A Fourier analysis shows that the IFT derived time domain modulation is intrinsically rippled at the frame scale, which can suppress attention to key frames in low gain intervals and induce sharp changes between adjacent frames.
@@
    We introduce Phase Aggregated Smoothing (PAS), a plug and play, training-free mechanism that assigns opposed phase offsets to different heads while shifting only the query stream. Each head preserves the original frequency spectrum magnitude, and standard head aggregation yields a smoother effective modulation over time.
@@ 2 Related Work
Videos as Input to Transformers. Early Vision Transformers for video showed that attention-only backbones can rival convolutional designs when space and time attention as well as tokenization are carefully structured [dosovitskiy2021imageworth16x16words]. TimeSformer decomposes attention along spatial and temporal axes to control cost [bertasius2021spacetimeattentionneedvideo]. ViViT popularizes tubelet embedding by merging patches across adjacent frames so that each token summarizes a short temporal slice (a “bin”) [arnab2021vivitvideovisiontransformer]. Hierarchical variants such as Video Swin adopt shifted local windows for better efficiency [liu2021videoswintransformer]. Multiscale families, for example MViT and memory-augmented successors, trade resolution and depth over time to handle long clips within fixed budgets [fan2021multiscalevisiontransformers]. These designs motivate low frame rates and frame merging in practice, which in turn increases temporal phase sensitivity. Our approach is orthogonal to these architectures and aims to improve robustness to sampling offsets without changing token budgets or backbones.
@@
Positional Encodings for Video and Multimodal Settings. Rotary positional embeddings (RoPE) inject position via complex-plane rotations and are widely used in language and vision Transformers [su2023roformerenhancedtransformerrotary]. Extending RoPE to images and videos, including M-RoPE, raises questions about axis coupling and frequency allocation. This motivated vision-specific analyses of 2D and 3D rotary designs, as well as multimodal formulations that interleave temporal and spatial channels in a unified scheme [heo2024rotarypositionembeddingvision, wei2025videoropemakesgoodvideo, wang2024qwen2vlenhancingvisionlanguagemodels]. Alternatives based on relative positions or lightweight gating offer different trade-offs between expressiveness and cost [Hao_2024]. Our work focuses on absolute temporal encoding and studies temporal robustness from a Fourier perspective.
@@
Temporal Robustness: Frame Rates, Sampling Offsets, and Test-time Strategies. A substantial body of work shows that action recognition accuracy is sensitive to frame rate and sampling choices. Lowering FPS or shifting sampling windows can change which motion cues are aggregated per token and may flip predictions on borderline cases [schiappa2023largescalerobustnessanalysisvideo]. Benchmarks that emphasize short-term object interactions further expose this fragility under sparse sampling [10.1145/2425836.2425909]. Architectural responses include dual-rate pathways that separate slow semantic context from fast motion dynamics [Feichtenhofer_2019_ICCV]. At inference time, multi-pass strategies improve stability by averaging predictions across resampled offsets or by temporally consistent adaptation, but they increase latency and complexity [yi2023temporalcoherenttesttimeoptimization]. In contrast, we target single-pass coverage of multiple temporal phases by distributing small phase offsets across attention heads.
@@
Training-free baselines. Complementary to the efficiency and positional-encoding lines above, several training-free Video-LLM strategies expand temporal coverage at inference. SlowFast-LLaVA constructs dual pathways, where the Slow branch uses low FPS and high spatial detail and the Fast branch uses high FPS with aggressive downsampling. The two streams are concatenated to stabilize decisions under sparse sampling and motion aliasing [xu2024slowfastllavastrongtrainingfreebaseline]. TS-LLaVA (“thumbnail and sampling”) mixes a compact thumbnail that preserves global spatial cues with a lightly sampled temporal stream, providing broader time coverage without changing weights [qu2024tsllavaconstructingvisualtokens]. They can be layered on top of the same backbone as ours and directly address failure modes induced by low FPS and frame merging, which makes them ideal comparators to our approach.
@@ 3.4 Implementation
Engineering details. (i) Hook point: apply $\Gamma_{\delta}$ after the base positional encoding is applied to $Q$. (ii) Masking: ensure only video tokens are affected. (iii) MHA [vaswani2023attentionneed] / GQA [ainslie2023gqatraininggeneralizedmultiquery] compatibility: broadcast offsets along the per-head dimensions of $Q$. (iv) Unit conversion: keep consistency with the time definition when using bin or frame units.
@@ 4 Experiment
4.1 Datasets and Evaluation
Action recognition. 20BN-Jester [9022297] emphasizes directional and ordering cues. Something-Something V2 [Goyal_2017_ICCV] focuses on object manipulations, and Kinetics-700 [smaira2020shortnotekinetics7002020human] covers diverse motions. For these three datasets, the sampling satisfies Nyquist, so PAS is tested in a phase-stable regime. Breakfast Actions [Kuehne_2014_CVPR] contains long activities; we use sub-sampled clips that are below Nyquist to assess behavior under aliasing.
General video-LLM suites. MVBench [li2024mvbenchcomprehensivemultimodalvideo] targets quasi static perception, TempCompass [liu2024tempcompassvideollmsreally] isolates temporal perception, PerceptionTest [pătrăucean2023perceptiontestdiagnosticbenchmark] probes multimodal reasoning, EgoSchema [mangalam2023egoschemadiagnosticbenchmarklongform] evaluates long horizon understanding, and MMBench-Video [fang2024mmbenchvideolongformmultishotbenchmark] measures multi shot and long form comprehension.
Table 1: Benchmarks used in our evaluation. The Nyquist column indicates whether the effective temporal sampling meets the Nyquist condition for the temporal RoPE line set. Only Breakfast is sub-Nyquist.
Dataset Type Key challenge Nyquist?
20BN-Jester [9022297] Action Directional cues $\checkmark$
Something-Something V2 [Goyal_2017_ICCV] Action Object manipulation $\checkmark$
Kinetics-700 [smaira2020shortnotekinetics7002020human] Action Diverse motions $\checkmark$
Breakfast Actions [Kuehne_2014_CVPR] Action Long, sub-sampled $\times$
MVBench [li2024mvbenchcomprehensivemultimodalvideo] General Motion analysis $\checkmark$
TempCompass [liu2024tempcompassvideollmsreally] General Temporal dependency $\checkmark$
EgoSchema [mangalam2023egoschemadiagnosticbenchmarklongform] General Long horizon $\checkmark$
MMBench-Video [fang2024mmbenchvideolongformmultishotbenchmark] General Multi shot $\checkmark$
4.2 Main Results
@@
SlowFast-LLaVA [xu2024slowfastllavastrongtrainingfreebaseline] 14.9 9.3 16.8 14.7 45.1 42.3 44.2 41.7 69.2 73.5 68.0 63.6 1.76
TS-LLaVA [qu2024tsllavaconstructingvisualtokens] 15.4 8.9 15.1 12.6 44.7 40.8 43.6 41.0 67.8 71.4 69.3 65.8 1.70
@@ 4.4 Ablation: Impact of Sampling Rate on PAS
Setup. All settings match section 4.2 except that $r$ is the independent variable. For each $r$, we compare the original backbone and PAS under identical preprocessing and decoding, with the same token budget at each sampling ratio $r$. PAS uses K=2K{=}2 head groups with per-head offsets $(0,\Delta)$ fixed at $\Delta{=}0.5$ (bin units), applied only to the query stream after the base positional encoding. We report accuracy (Acc) for all four datasets.

### v2

@@ Abstract
Video LLMs suffer from temporal inconsistency: small shifts in frame timing can flip attention and suppress relevant frames. We trace this instability to the common extension of Rotary Position Embeddings to video through multimodal RoPE. The induced inverse Fourier time kernel exhibits frame-scale ripples that multiply adjacent frames by different factors, which perturbs attention that should otherwise be governed by the raw query key inner product. We present Phase Aggregated Smoothing (PAS), a simple, training-free mechanism that applies small opposed phase offsets across heads and then aggregates their outputs. PAS preserves the per-head spectrum magnitude, while the aggregation effectively smooths the temporal kernel and reduces phase sensitivity without changing the positional encoding structure. Our analysis shows that the RoPE rotated logit can be approximated as a content dot product scaled by a time kernel; smoothing this kernel yields Lipschitz stability of attention to small temporal shifts; multi-phase averaging attenuates high frequency ripples while preserving per-head spectra under Nyquist-valid sampling. Experiments on multiple video understanding benchmarks under matched token budgets show consistent improvements with negligible computational overhead. PAS provides a plug-and-play upgrade for robust temporal encoding in Video LLMs.
@@ 1 Introduction
Large multimodal language models extend language-centric architectures to video, enabling open-ended description and reasoning over long temporal contexts. A widely used positional encoding is Rotary Position Embeddings (RoPE), originally designed for discrete text tokens. In Video LLMs, RoPE is generalized to a multimodal variant (M-RoPE) that encodes temporal ($T$), height ($H$), and width ($W$) dimensions. Along the temporal axis, this construction amounts to assigning a finite set of temporal frequencies; each frequency line rotates query and key pairs and induces an independent, periodic contribution to the attention logit. In practice, downstream accuracy is sensitive to sampling details: small changes of frame rate or sampling offsets can alter which motion cues are aggregated per token, exposing fragility in temporal encoding. One of the reasons is that when this text-oriented mechanism is applied to continuous time, the attention distribution may be disrupted.
Figure 1: A failure from a real clip [2]. Temporal RoPE imposes an interval-dependent gain on attention; when a key frame lies in a low-gain trough of the modulation, it is down weighted compared to less relevant frames, which propagates to downstream errors.
@@
We adopt a Fourier view to address this issue. Let $\{\omega_{k}\}$ denote the temporal line set and define $m(t)$ as the inverse Fourier time-domain kernel associated with these lines. Under a standard high dimensional approximation, the RoPE rotated logit at relative lag $t$ can be written as the unrotated content dot product multiplied by $\mathrm{Re}\{m(t)\}$; we later formalize this phase modulation view as Theorem 1. Because $m(t)$ is an average of cosines over $\{\omega_{k}\}$, it is intrinsically rippled at the frame scale, which renders the overall modulation non-smooth. Such non-smoothness means that the attention assigned to frames varies with their relative lag: within specific temporal intervals, the modulation can either amplify or attenuate attention. A concrete failure occurs when a key frame lands in a low gain interval of $m(t)$, causing its evidence to be suppressed or even ignored (Figure 1). The upper panel of Figure 2 visualizes how this interval-dependent modulation undermines temporal stability by imposing gain swings across adjacent frames.
@@
    We identify and formalize a core limitation of applying text-oriented RoPE to the time axis of video. A Fourier analysis shows that the IFT derived time domain modulation is intrinsically rippled at the frame scale, which can suppress attention to key frames in low gain intervals and induce sharp changes between adjacent frames.
@@
    We introduce Phase Aggregated Smoothing (PAS), a plug-and-play, training-free mechanism that assigns opposed phase offsets to different heads while shifting only the query stream. Each head preserves the original frequency spectrum magnitude, and standard head aggregation yields a smoother effective modulation over time.
@@ 2 Related Work
Videos as Input to Transformers. Early Vision Transformers (e.g., TimeSformer [4], ViViT [3], Video Swin [16], and MViT [6]) structure space-time attention and tokenization to control computational cost [5]. These designs motivate low frame rates and frame merging in practice, which in turn increases temporal phase sensitivity. Our approach is orthogonal to these architectures and aims to improve robustness to sampling offsets without changing token budgets or backbones.
@@
Positional Encodings for Video and Multimodal Settings. Rotary positional embeddings (RoPE) inject position via complex-plane rotations and are widely used in language and vision Transformers [22]. Extending RoPE to images and videos, including M-RoPE, raises questions about axis coupling and frequency allocation. This motivated vision-specific analyses of 2D and 3D rotary designs, as well as multimodal formulations that interleave temporal and spatial channels in a unified scheme [12, 25, 24]. Alternatives based on relative positions or lightweight gating offer different trade-offs between expressiveness and cost [10]. Our work focuses on absolute temporal encoding and studies temporal robustness from a Fourier perspective.
@@
Temporal Robustness: Frame Rates, Sampling Offsets, and Test-time Strategies. A substantial body of work shows that action recognition accuracy is sensitive to frame rate and sampling choices. Lowering FPS or shifting sampling windows can change which motion cues are aggregated per token and may flip predictions on borderline cases [20]. Benchmarks that emphasize short-term object interactions further expose this fragility under sparse sampling [11]. Architectural responses include dual-rate pathways that separate slow semantic context from fast motion dynamics [8]. At inference time, multi-pass strategies improve stability by averaging predictions across resampled offsets or by temporally consistent adaptation, but they increase latency and complexity [27]. In contrast, we target single-pass coverage of multiple temporal phases by distributing small phase offsets across attention heads.
@@
Training-free baselines. Complementary to the efficiency and positional-encoding lines above, several training-free Video-LLM strategies expand temporal coverage at inference. SlowFast-LLaVA constructs dual pathways, where the Slow branch uses low FPS and high spatial detail and the Fast branch uses high FPS with aggressive downsampling. The two streams are concatenated to stabilize decisions under sparse sampling and motion aliasing [26]. TS-LLaVA (“thumbnail and sampling”) mixes a compact thumbnail that preserves global spatial cues with a lightly sampled temporal stream, providing broader time coverage without changing weights [19]. They can be layered on top of the same backbone as ours and directly address failure modes induced by low FPS and frame merging, which makes them ideal comparators to our approach.
@@ 3.4 Implementation
Engineering details. (i) Hook point: apply $\Gamma_{\delta}$ after the base positional encoding is applied to $Q$. (ii) Masking: ensure only video tokens are affected. (iii) MHA [23] / GQA [1] compatibility: broadcast offsets along the per-head dimensions of $Q$. (iv) Unit conversion: keep consistency with the time definition when using bin or frame units.
@@ 4 Experiment
SlowFast-LLaVA [26] 14.9 9.3 16.8 14.7 45.1 42.3 44.2 41.7 69.2 73.5 68.0 63.6 1.76
TS-LLaVA [19] 15.4 8.9 15.1 12.6 44.7 40.8 43.6 41.0 67.8 71.4 69.3 65.8 1.70
@@
4.1 Datasets and Evaluation
Action recognition. 20BN-Jester [18] emphasizes directional and ordering cues. Something-Something V2 [9] focuses on object manipulations, and Kinetics-700 [21] covers diverse motions. For these three datasets, the sampling satisfies Nyquist, so PAS is tested in a phase-stable regime. Breakfast Actions [13] contains long activities; we use sub-sampled clips that are below Nyquist to assess behavior under aliasing.
General video-LLM suites. MVBench [14] targets quasi-static perception, TempCompass [15] isolates temporal perception, PerceptionTest [pătrăucean2023perceptiontestdiagnosticbenchmark] probes multimodal reasoning, EgoSchema [17] evaluates long-horizon understanding, and MMBench-Video [7] measures multi-shot comprehension.
Table 2: Benchmarks used in our evaluation. The Nyquist column indicates whether the effective temporal sampling meets the Nyquist condition for the temporal RoPE line set.
Dataset Type Key challenge Nyquist?
20BN-Jester [18] Action Directional cues $\checkmark$
Something-Something V2 [9] Action Object manipulation $\checkmark$
Kinetics-700 [21] Action Diverse motions $\checkmark$
Breakfast Actions [13] Action Long, sub-sampled $\times$
MVBench [14] General Motion analysis $\checkmark$
TempCompass [15] General Temporal dependency $\checkmark$
EgoSchema [17] General Long-horizon $\checkmark$
MMBench-Video [7] General Multi-shot $\checkmark$
4.2 Main Results
@@ 4.4 Ablation: Impact of Sampling Rate on PAS
Setup. All settings match Section 4.2 except that $r$ is the independent variable. For each $r$, we compare the original backbone and PAS under identical preprocessing and decoding, with the same token budget at each sampling ratio $r$. PAS uses K=2K{=}2 head groups with per-head offsets $(0,\Delta)$ fixed at $\Delta{=}0.5$ (bin units), applied only to the query stream after the base positional encoding.